### v1

# Doublons Bloch oscillations in the mass-imbalanced extended Fermi-Hubbard model

Abstract
Interactions between particles normally induce the decay of the particles Bloch oscillations (BOs) in a periodic lattice. In the large on-site interactions region, spin-$1/2$ fermions may form into doublon bound state and undergoes doublon BOs in the present of tilted potential. Here we investigate the impact of nearest-neighbor interaction $V$ on the multi-doublon BOs in a mass-imbalanced extended Fermi-Hubbard model. We derive a doublons effective Hamiltonian, and show that a minor change of $V$ can alter the dynamic behaviors of doublons qualitatively. Notably, at a resonance point, the doublons act like free hard-core bosons. Under a tilted potential, the system may exhibit different types of multi-doublon BOs at or deviation from the resonance point with different initial states. Numerical results are presented to demonstrate our conclusions in the one- and two-dimensional systems.
I Introduction
Bloch oscillations (BOs) of electrons in crystal lattices were first predicted by Felix Bloch in 1929 [1]. When a finite dc electric field is applied to the ideal crystal lattice, the eigenstates of electrons become localized, accompanied by the ladder-like energy levels, i.e., the Wannier-Stark (WS) ladder [2], resulting in the coherent motion of wave packet, which is the underlying mechanism of the BOs. Until sixty years after being predicted, the BOs were experimentally observed in semiconductor superlattices [3, 4] and optical lattices [5, 6, 7, 8]. In recent years, the BOs have been intensively explored in strongly correlated systems, including quantum spin chain [9, 10, 11, 12, 13, 14, 15, 16], interacting boson and fermion systems [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32], in both theoretical and experimental aspects.
In general, interactions between particles induce the decoherence of BOs [17, 18, 19, 20, 21, 22]. Nevertheless, a strong interaction strength may bound two particles into a doublon [33, 34, 35, 29, 30, 36, 37, 38, 39, 40, 41, 42] or a bound pair [43, 44, 45, 46, 27, 47, 48, 49] as high-energy excitations, in the boson and fermion systems, respectively. The BOs of bound particles in these systems exhibit characteristics different from the free-particle situations, for example, the fractional BOs [27, 50] or frequency doubling phenomena [23, 24, 25, 51] have been observed. The Fermi-Hubbard model [52, 53] describes a typical strongly correlated electronic system, with hopping amplitude $J$ and on-site interaction with strength $U$. A more general consideration leads to the extended Fermi-Hubbard model [29, 36, 54, 55, 56, 57, 58], which includes the nearest-neighbor (NN) interaction with strength $V$. Obviously, both the on-site and NN interactions can affect the behavior of many-body BOs. The aforementioned research works have studied the BOs in the (extended) Fermi-Hubbard chains, but mainly focused on the one doublon or two fermions case. It is natural and interesting to take the doublon-doublon interaction into account in the dynamics of doublons BOs.
This paper aims to investigate the doublons BOs in the extended Hubbard model with imbalanced mass of spin-up and spin-down fermions [37, 59, 60, 61, 62, 63, 64, 65, 66, 67], under a tilted potential. We derive a doublons effective Hamiltonian in the strong on-site interaction situation to capture the physics of the doublons dynamics. We show that the NN interaction induces the bound-doublon eigenstates, which dominate the doublons dynamics. It is observed that (i) at a resonance point, the bound-doublon eigenstates disappear, and the doublons act like free hard-core bosons; (ii) when the NN interaction is strong enough, the energy levels of bound doublons are completely isolated from the scattering band of doublons. These properties enable the doublons BOs to occur after applying the tilted potential for the systems in the above two cases. We present the numerical simulations of time evolutions to verify our analysis. For the above two cases, the breathing modes of BOs exhibit the same pattern of doublon density, but with significant difference of mean-square displacements (MSD) of doublons in one dimension ($1$D). The numerical results of doublons BOs for the system on a square lattice are also presented and discussed. Our results show that the multi-doublon BOs can survive in spite of the present of the NN interaction of fermions.
The rest of this paper is organized as follows: In Sec. II, we introduce the Hamiltonian of mass-imbalanced extended Fermi-Hubbard model, and derive its doublons effective Hamiltonian. In Sec. III, we analyze the mechanisms of the two types of doublons BOs, with the numerical results as the verifications. In Sec. IV, we discuss the doublons dynamics of the system on a square lattice. Finally, we conclude and discuss our findings in Sec. V.
II Model and doublons effective Hamiltonian
The Hamiltonian of the system we considered can be written as
[…]
where
$\hat{H}$ $=$ $-\sum_{\left\langle i,j\right\rangle}\sum_{\sigma=\uparrow,\downarrow}J_{\sigma}\hat{c}_{i,\sigma}^{\dagger}\hat{c}_{j,\sigma}+\mathrm{H.c.}$ (2)
$+U\sum_{i}\hat{n}_{i,\uparrow}\hat{n}_{i,\downarrow}+\frac{1}{2}V\sum_{\left\langle i,j\right\rangle}\hat{n}_{i}\hat{n}_{j},$
is the extended Fermi-Hubbard model with imbalance mass. Figure 1 presents a schematic illustration of the system in $1$D. For a $1$D system, the dc field term with strength $F$ takes the form of tilted potential
$\hat{H}_{\mathrm{dc}}=F\sum_{j}j\hat{n}_{j},$ (3)
which is readily feasible in the experiments of ultracold atoms [68, 22, 69]. Here, $\hat{c}_{i,\sigma}^{\dagger}$ ($\hat{c}_{i,\sigma}$) is the creation (annihilation) operator of fermion with spin $\sigma$ on the site with index $i$; the summation $\sum_{\left\langle i,j\right\rangle}$ is over the NN sites with indices $i$ and $j$; and $\hat{n}_{i,\sigma}=\hat{c}_{i,\sigma}^{\dagger}\hat{c}_{i,\sigma}$, $\hat{n}_{i}=\hat{n}_{i,\uparrow}+\hat{n}_{i,\downarrow}$ are the corresponding fermion number operators. The strength of on-site interaction and NN interaction are $U$ and $V$, respectively, which can be positive (repulsion) or negative (attraction). The mass of the spin-up and spin-down fermions are different, manifesting as the difference of the corresponding hopping amplitudes. This has been realized with Fermi gas of ${}^{171}\mathrm{Yb}$ atoms in a state-dependent optical lattice recently [65]. We assume $J_{\downarrow}>J_{\uparrow}>0$ in the following discussions. It has been shown that the heavy particles acting as an effective disorder potential for the light ones in a extremely strong mass imbalance [59]. The system conserves the total fermion number $\hat{n}=\sum_{i}\hat{n}_{i}$, as well as the total spin polarization $\hat{s}^{z}=\sum_{i}(\hat{n}_{i,\uparrow}-\hat{n}_{i,\downarrow})/2$.
Figure 1: Schematic illustration of the mass-imbalanced extended Fermi-Hubbard chain with the tilted potential $F$. Here the filled yellow polygons and gray ellipse represent the on-site interaction $U$ and NN interaction $V$, respectively. The hopping amplitudes for the heavy (red) and light (blue) fermions are $J_{\uparrow}$ and $J_{\downarrow}$ ($J_{\downarrow}>J_{\uparrow}>0$), respectively.
The standard Hubbard model in 1D was exactly solved by Lieb and Wu with the Bethe ansatz method [70]. The imbalance of mass $J_{\uparrow}\neq J_{\downarrow}$ and nonzero NN interaction $V\neq 0$ resists the exact analytical treatments so far. Nevertheless, we pursue the perturbation and numerical solutions. In this work, we focus on the dynamics of the system with multidoublon, i.e., each site of the system is either doubly occupied or empty. This invariant subspace, referred to as doublons subspace, is approximately achieved when considering the strong on-site interaction region $\left|U\right|\gg J_{\sigma},\left|V\right|,\left|F\right|$. In the following, we derive the effective Hamiltonian of this case.
To explore the feasibility of the doublons BOs in the present model, we shall derive the doublons effective Hamiltonian. We firstly focus on the Hamiltonian without the tilted potential. In the strong on-site interaction region, the states in the doublons subspace are well separated from others in the spectrum. The doublons effective Hamiltonian for $\hat{H}$ can be obtained perturbatively through the Schrieffer-Wolff transformation [71, 72, 73, 74, 75]
$\hat{H}_{\mathrm{eff}}$ $=$ $e^{\mathrm{i}\hat{S}}\hat{H}e^{-\mathrm{i}\hat{S}}$ (4)
$=$ $\hat{H}+\mathrm{i}\left[\hat{S},\hat{H}\right]+\frac{\mathrm{i}^{2}}{2}\left[\hat{S},\left[\hat{S},\hat{H}\right]\right]+...\ ,$
where $\mathrm{i}=\sqrt{-1}$. The terms that change the number of doublons can be eliminated by properly selecting the Hermitian operator $\hat{S}$. Up to second order, the doublons effective Hamiltonian reads (see the APPENDIX for detailed calculation)
$\hat{H}_{\mathrm{eff}}=\sum_{\left\langle i,j\right\rangle}\left[J_{\mathrm{eff}}\left(\hat{d}_{i}^{\dagger}\hat{d}_{j}+\hat{d}_{j}^{\dagger}\hat{d}_{i}\right)+\frac{1}{2}V_{\mathrm{eff}}\hat{n}_{i}\hat{n}_{j}\right],$ (5)
where $d_{i}^{\dagger}=\hat{c}_{i,\uparrow}^{\dagger}\hat{c}_{i,\downarrow}^{\dagger}$ is the creation operator of doublon; the effective hopping amplitude and NN interaction of doublons are $J_{\mathrm{eff}}=2J_{\uparrow}J_{\downarrow}/U$ and $V_{\mathrm{eff}}=V-(J_{\uparrow}^{2}+J_{\downarrow}^{2})/U$, respectively. This means that the doublons move in the lattice through the second-order processes, and all the three parameters of the Hubbard model play a crucial part in the doublon-doublon interaction. Clearly, the doublons satisfies the statistics of hard-core bosons, that is
$\left\{\hat{d}_{i},\hat{d}_{i}^{\dagger}\right\}$ $=$ $1,\left\{\hat{d}_{i},\hat{d}_{i}\right\}=0,$
$\left[\hat{d}_{i},\hat{d}_{j}^{\dagger}\right]$ $=$ $\left[\hat{d}_{i},\hat{d}_{j}\right]=0,(i\neq j).$ (6)
That is to say, $\hat{H}_{\mathrm{eff}}$ in Eq. (5) suggests that in the strong on-site interaction region, the nonequilibrium dynamics of doublons in the mass-imbalanced extended Hubbard model Eq. (2) resembles that in a hard-core bosons system with effective hopping amplitude $J_{\mathrm{eff}}$ and NN interaction $V_{\mathrm{eff}}/2$. This is distinguished from the half-filled ground state in the case of $U>0$, the effective Hamiltonian of which is the spin-$1/2$ anisotropic XXZ Heisenberg chain [60], and the introducing of NN interaction $V$ only gives a constant term in that. Notably, the above result is independent of spatial dimension.
III Doublons Bloch oscillations
We have seen that the NN interaction $V$ plays a vital role in the doublon-doublon interaction from the doublons effective Hamiltonian Eq. (5). It is worth noting that when the energies of interactions and hoppings resonate, the doublon-doublon interaction vanishes. In this section, we investigate the doublons BOs under the tilted potential $F$ in two parameter regions.
III.1 Two-doublon Bloch oscillations
In order to gain insight into the doublon dynamics under the tilted potential, we first analyze the two-doublon situation in a chain with length $L$. Due to the translational invariance of the lattice, the effective Hamiltonian Eq. (5) in two-doublon subspace can be written as
$\hat{H}_{\mathrm{eff}}^{(2)}(K)$ $=$ $2J_{\mathrm{eff}}\cos\left(\frac{K}{2}\right)\sum_{r=1}^{L-1}\left[\left|\phi_{r}(K)\right\rangle\left\langle\phi_{r+1}(K)\right|+\mathrm{H.c.}\right]$ (7)
$+\frac{1}{2}V_{\mathrm{eff}}\left|\phi_{1}(K)\right\rangle\left\langle\phi_{1}(K)\right|,$
under the two-doublon basis with momentum $K$
$\left|\phi_{r}(K)\right\rangle=\frac{1}{\sqrt{L}}e^{\mathrm{i}Kr/2}\sum_{j}e^{\mathrm{i}Kj}\hat{d}_{j}^{\dagger}\hat{d}_{j+r}^{\dagger}\left|\text{{Vac}}\right\rangle,$ (8)
where the vacuum state is defined as $\hat{d}_{i}\left|\text{{Vac}}\right\rangle=0$, and $r=1,2,3,...,L-1$ is the distance between two doublons. Note that $\hat{H}_{\mathrm{eff}}^{(2)}(K)$ can be regarded as the single-particle Hamiltonian of a tight-binding chain with hopping amplitude $2J_{\mathrm{eff}}\cos(K/2)$, and with impurity potential $V_{\mathrm{eff}}/2$ at one end. Such impurity model supports bound state in certain parameter region. In order to solve the eigenvalue equation $\hat{H}_{\mathrm{eff}}^{(2)}(K)\left|\psi(K)\right\rangle=E(K)\left|\psi(K)\right\rangle$, we expand the eigenstate $\left|\psi(K)\right\rangle$ as
$\left|\psi(K)\right\rangle=\sum_{r=1}^{L-1}C_{r}(K)\left|\phi_{r}(K)\right\rangle.$ (9)
Substituting this into the eigenvalue equation, we obtain
$2J_{\mathrm{eff}}\cos\left(\frac{K}{2}\right)\left[C_{r+1}(K)+C_{r-1}(K)\right]$
$+\left[\frac{1}{2}V_{\mathrm{eff}}\delta_{r,1}-E(K)\right]C_{r}(K)=0,$ (10)
with boundary condition being $C_{0}(K)=0$. All the eigenstates and eigenenergies of Hamiltonian Eq. (7) can be obtained by solving this equation, numerically or analytically.
Next we focus on the bound-doublon solution. We takes the ansatz for the coefficient of the bound state as $C_{r}(K)=\exp[-\beta(K)r]$ with $\beta(K)>0$ being the decay rate of wave function. Substituting it into Eq. (10), straightforward calculation gives the decay rate
$\beta(K)=\ln\left[\frac{V_{\mathrm{eff}}}{4J_{\mathrm{eff}}\cos\left(K/2\right)}\right],$ (11)
and the bound-doublon energy
$E_{\textrm{bou.}}(K)=\frac{1}{2}V_{\mathrm{eff}}+\frac{8J_{\mathrm{eff}}^{2}\cos^{2}\left(K/2\right)}{V_{\mathrm{eff}}}.$ (12)
The bound-state condition $\beta(K)>0$ requires $\left|V_{\mathrm{eff}}\right|>4\left|J_{\mathrm{eff}}\cos\left(K/2\right)\right|$. In this case, the doublons bound states are induced by the NN interaction $V$. In Fig. 2, we present the two-doublon spectrum for various NN interaction strengths, which can be obtained by diagonalizing the two-doublon effective Hamiltonian Eq. (7) or solving Eq. (10). In Fig. 2(a), we can see that when $V_{\mathrm{eff}}=0$, the spectrum is all composed of scattered band of doublons. Increasing $\left|V\right|$ would induce the isolated energy levels, which are the energies of bound doublons [see Fig. 2(b)]. When $\left|V_{\mathrm{eff}}\right|>4\left|J_{\mathrm{eff}}\cos\left(K/2\right)\right|$ for all $K$, the bound-doublon band is completely isolated from the scattered band [see Fig. 2(c)]. In the following, we shall investigate the BOs in the parameter regions of Figs. 2(c) and (a) sequentially.
Figure 2: Two-doublon spectrum as a function of momentum $K$ for various NN interactions, which are obtained from the exact diagonalization of the two-doublon effective Hamiltonian in Eq. (7). Parameters of the systems are (a) $V=-1/6$, (b) $V=-0.8$ and (c) $V=-2$. Other parameters are $L=100$, $U=-6$, $J_{\downarrow}=1$ and $J_{\uparrow}=0.5$. The doublons scattered bands and bound-doublon energy levels are indicated by the black and red lines, respectively. Here the energy is in unit of $J_{\downarrow}$.
Here we first focus on the dynamics of the two bound doublons. For a large NN interaction $V$, we have $\left|V_{\mathrm{eff}}\right|\gg 4\left|J_{\mathrm{eff}}\cos\left(K/2\right)\right|$. In this case, two doublons in the bound band are strongly bound. Then we are able to write down the two-doublon effective Hamiltonian under the strongly bound basis $\left|\varphi_{j}\right\rangle=\hat{d}_{j}^{\dagger}\hat{d}_{j+1}^{\dagger}\left|\text{{Vac}}\right\rangle$. We divide the effective Hamiltonian Eq. (5) into two parts, i.e., $\hat{H}_{\mathrm{eff}}^{(0)}=V_{\mathrm{eff}}\hat{n}_{i}\hat{n}_{j}/2$ and $\hat{H}_{\mathrm{eff}}^{\prime}=J_{\mathrm{eff}}(\hat{d}_{i}^{\dagger}\hat{d}_{j}+\hat{d}_{j}^{\dagger}\hat{d}_{i})$, and treat $\hat{H}_{\mathrm{eff}}^{\prime}$ as a perturbation. According to the degenerate perturbation theory, the effective Hamiltonian is given by the following formula
$\hat{H}_{\mathrm{eff}}^{(2,\mathrm{bou.})}$ $=$ $\hat{P}\hat{H}_{\mathrm{eff}}^{(0)}\hat{P}+\hat{P}\hat{H}_{\mathrm{eff}}^{\prime}\hat{Q}\frac{1}{E_{\mathrm{eff}}^{(0)}-\hat{H}_{\mathrm{eff}}^{(0)}}\hat{Q}\hat{H}_{\mathrm{eff}}^{\prime}\hat{P}$ (13)
$+O\left(J_{\mathrm{eff}}^{3}/V_{\mathrm{eff}}^{2}\right),$
up to second order $J_{\mathrm{eff}}^{2}/V_{\mathrm{eff}}$. Here the projectors are defined as $\hat{P}=\sum_{j}\left|\varphi_{j}\right\rangle\left\langle\varphi_{j}\right|$ and $\hat{Q}=\hat{1}-\hat{P}$, and the energy of the unperturbed state is $E_{\mathrm{eff}}^{(0)}=V_{\mathrm{eff}}/2$. Direct calculation shows that
$\hat{H}_{\mathrm{eff}}^{(2,\mathrm{bou.})}\approx\frac{2J_{\mathrm{eff}}^{2}}{V_{\mathrm{eff}}}\sum_{j}\left(\left|\varphi_{j}\right\rangle\left\langle\varphi_{j+1}\right|+\left|\varphi_{j+1}\right\rangle\left\langle\varphi_{j}\right|\right),$ (14)
in which the constant potential that is proportional to $\sum_{j}\left|\varphi_{j}\right\rangle\left\langle\varphi_{j}\right|$ is omitted. Now the bound-doublon physics is clear. From Eq. (14), we can see that the bound-doublon hops in the lattice through a second-order process of the single-doublon hopping.
Figure 3: Numerical results of the time evolution of doublon density defined in Eq. (17) for different initial states with (a) delta distribution in Eq. (15) and (b) Gaussian distribution in Eq. (16). Parameters of the system are set as the same as that in Fig. 2 (c), except that the system size is $L=30$. The strength of tilted potential is taken as $F=0.001$. Parameters of the Gaussian wave packet are $\alpha=0.4$, $j_{0}=15$ and $K_{0}=\pi/2$. The time unit is $J_{\downarrow}^{-1}$, and we set $J_{\downarrow}=1$ throughout the numerical simulations in this work.
When the tilted potential $\hat{H}_{\mathrm{dc}}$ in Eq. (3) is applied, the quadrupled dc field $4F\sum_{j}j\left|\varphi_{j}\right\rangle\left\langle\varphi_{j}\right|$ arises in the bound-doublon subspace, forming the WS ladder for bound doublons. Then the bound doublons may undergo BOs. The oscillating period would be $T=2\pi/(4F)$, which is a quarter of that of a free particle in lattice [76]. To verify the above analysis, we conduct numerical simulations of time evolutions. We consider the delta and Gaussian distributions of initial bound-doublon states, which are
$\left|\Psi_{\mathrm{a}}\left(0\right)\right>=\hat{d}_{15}^{\dagger}\hat{d}_{16}^{\dagger}\left|\text{{Vac}}\right\rangle,$ (15)
and
$\left|\Psi_{\mathrm{b}}\left(0\right)\right>=\Omega^{-1}\sum_{j}e^{-\alpha^{2}(j-j_{0})^{2}+\mathrm{i}K_{0}j}\hat{d}_{j}^{\dagger}\hat{d}_{j+1}^{\dagger}\left|\text{{Vac}}\right\rangle,$ (16)
[…]
$\hat{n}_{j,\uparrow}\hat{n}_{j,\downarrow}\right>=\left\langle\Psi\left(t\right)\right|\hat{n}_{j,\uparrow}\hat{n}_{j,\downarrow}\left|\Psi\left(t\right)\right>.$ (17)
The numerical results are presented in Fig. 3, which are consistent with our analysis. Figures. 3(a) and (b) show that two different initial excitations undergo the breathing mode and oscillating mode, respectively. The oscillating periods are both $T=2\pi/(4F)\approx 1570.80$.
Figure 4: Numerical results of the time evolution of doublon density for different NN interaction $V$, tilted potential $F$ and doublon number. The system sizes for the two-doublon cases in (a)-(d) and four-doublon cases in (e)-(f) are taken as $L=30$ and $L=20$, respectively. Other system parameters are $U=-10$, $J_{\downarrow}=1$ and $J_{\uparrow}=0.1$.
Figure 5: Mean-square displacement between the doublons [see Eq. (18)] as a function of time for the cases in Fig. 3(a) and Figs. 4(a)-(d). The time scales are normalized by the final time of the data in each case.
III.2 Free doublons dynamics in an interacting system
Now we investigate the BOs at the resonance point of system parameters, which is $V_{\mathrm{eff}}=V-(J_{\uparrow}^{2}+J_{\downarrow}^{2})/U=0$. We are interested in the multi-doublon dynamics. Obviously, the released doublons initial state acts like free hard-core bosons according to the doublons effective Hamiltonian in Eq. (5), but with renormalized NN hopping amplitude $J_{\mathrm{eff}}$. In Figs. 4(a) and (c), we present the numerical results of time evolution for the initial state in Eq. (15) with different NN interaction but without the tilted potential. It is shown that the deviation from the resonance condition $V=(J_{\uparrow}^{2}+J_{\downarrow}^{2})/U$ obstructs the transport of doublons. This can be demonstrated by the following correlation function that characterize the mean-square displacement (MSD) [78] between the doublons
$\overline{\mathcal{D}^{2}}(t)=\frac{1}{L}\sum_{i,j}(i-j)^{2}\left\langle\Psi\left(t\right)\right|\hat{n}_{i,\uparrow}\hat{n}_{i,\downarrow}\hat{n}_{j,\uparrow}\hat{n}_{j,\downarrow}\left|\Psi\left(t\right)\right>.$ (18)
The numerical results of the MSDs for these two cases are presented in Fig. 5. For the resonant case, we can observe the characteristic of ballistic transport, i.e., $\overline{\mathcal{D}^{2}}(t)\sim t^{2}$, which means the doublons spread without scattering.
When applying the tilted potential in Eq. (3), the doublons effective Hamiltonian becomes
[…]
where the doubling of the tilted potential comes form the double occupancy of fermions per site. This yields the WS ladder for free doublons. The period of the BOs would be a half of that of free particle in lattice. Numerical result in Fig. 4(d) shows the free-doublon BOs at the resonance point. From Figs. 4(a), (b) and the corresponding MSDs in Fig. 5, we can see that a minor deviation from the parameter $V=(J_{\uparrow}^{2}+J_{\downarrow}^{2})/U$ induces decoherence of the diffusion and BO, highlighting the effect of NN interaction. In Figs. 4(e) and (f), we present the results for the initial state with $4$ doublons ($8$ fermions)
$\left|\Psi\left(0\right)\right>=\prod_{i=9}^{12}\hat{d}_{i}^{\dagger}\left|\text{{Vac}}\right\rangle,$ (20)
corresponding to Figs. 4(c) and (d), but with a smaller system size $L=20$. The larger tilted potential in Fig. 4(f) is taken to constrain the amplitude of BO, avoiding the boundary effect. The diffusion behavior and BO are similar with the cases of $2$ doublons. The oscillating periods for the breathing modes are $T=2\pi/(2F)\approx 628.32$ and $314.16$ for the cases in Figs. 4(d) and (f), respectively.
We would like to stress that the breathing modes of BOs in Fig. 3(a) and Fig. 4(d) are different, although the profiles of their doublons densities look almost the same. Firstly, as we have pointed out, their oscillating periods are a quarter and a half of that of free particle in lattice, respectively. Secondly, in Fig. 3(a), two doublons are always bounded tightly together by the NN interaction, evolving as a bound-doublon wave packet, while in Fig. 4(d), the distance between two doublons changes periodically since they can be considered as free particles. The numerical results of the MSDs $\mathcal{D}(t)$ in Fig. 5 verify the above analysis. Nevertheless, one common point is that the spin-up and spin-down fermions are always well bounded as doublons, regardless of the significant difference in their mass.
IV Numerical results of two-dimensional system
So far, we have clarified the mechanism of doublons BOs in the $1$D mass-imbalanced extended Hubbard model. Since the doublons effective Hamiltonian is independent of spatial dimension, the generalization of doublons BOs to a higher dimension is straightforward for the resonance case $V=(J_{\uparrow}^{2}+J_{\downarrow}^{2})/U$. Here we perform numerical simulations of time evolutions for the $2$D system. The Hubbard model is taken as the Hamiltonian Eq. (2) on a square lattice with size $(L_{x},L_{y})=(10,9)$. While the $2$D tilted potential [79, 80] is
$\hat{H}_{\mathrm{dc}}^{(\mathrm{2D})}=\sum_{m,n}\mathbf{F}\cdot\mathbf{r}_{m,n}\hat{n}_{(m,n)},$ (21)
where $\mathbf{F}=(F_{x},F_{y})$ is the strength of $2$D tilted potential, and $\mathbf{r}_{m,n}=m\mathbf{x}+n\mathbf{y}$ is the lattice vector of the square lattice with unit vectors $\mathbf{x}$ and $\mathbf{y}$ along the $x$-axis and $y$-axis, respectively.
Taking two NN doublons localized in the center of the lattice as an initial state, Figs. 6(a) and (b) show the numerical results of doublon density at different time for zero and nonzero tilted potential $\mathbf{F}$, respectively. As expected, Fig. 6(a) shows that the doublons manifest the spreading dynamics resembling the free particles in a square lattice. While the doublons in Fig. 6(b) undergo the BOs, with periods $T_{x}=2\pi/(2F_{x})=104.72$ and $T_{y}=2\pi/(2F_{y})=52.36$ in $x$ and $y$ directions, respectively.
Figure 6: Snapshots of doublon density for the system on a square lattice at different time for (a) $F_{x}=F_{y}=0$, and (b) $F_{x}=0.03$, $F_{y}=0.06$. Other parameters of the system are $L_{x}=10$, $L_{y}=9$, $U=-10$, $V=(J_{\uparrow}^{2}+J_{\downarrow}^{2})/U=-0.101$, $J_{\downarrow}=1$ and $J_{\uparrow}=0.1$.
V Conclusion
We have studied multi-doublon BOs in the extended Fermi-Hubbard model with imbalance mass. The doublons effective Hamiltonian is given to elucidate the mechanisms of two types of doublons BOs, i.e., the bound-doublon and free-doublon BOs. The interaction between doublons stems not only from the NN interaction between fermions but also from the hopping and on-site interaction. Therefore, there is a resonance point in the system parameters that results in the free-doublon dynamics. It is shown that the spin-up and spin-down fermions are always well bounded as doublons and exhibit two types of BOs dynamics, despite of the imbalance of their mass. Numerical results of the time evolution of doublon density and MSD demonstrate our conclusions. The numerical simulations of the system on a square lattice suggest that the analysis can be applied to higher dimensions. Our findings highlight the role of NN interaction in the nonequilibrium dynamic properties, as well as pave the way for the future explorations of BOs in strongly correlated systems.
APPENDIX Derivation of the doublons effective Hamiltonian
Following the previous works on the $J/U$ perturbation expansion of the Fermi-Hubbard model using the Schrieffer-Wolff transformation [71, 72, 73, 74], we present a detailed derivation of the doublons effective Hamiltonian Eq. (5) in the strong on-site interaction region $\left|U\right|\gg J_{\sigma},\left|V\right|,\left|F\right|$. In this case, the states with sites being either doubly occupied or empty are well separated from others in the energy spectrum.
The Hamiltonian in Eq. (2) can be divided into five parts
$\hat{H}=\hat{H}_{J}^{+}+\hat{H}_{J}^{-}+\hat{H}_{J}^{0}+\hat{H}_{U}+\hat{H}_{V},$ (A1)
where $\hat{H}_{J}^{+}\ (\hat{H}_{J}^{-})$ describes the processes that increase (decrease) the number of doubly occupied sites by one (see Fig. A1), that are written as
$\hat{H}_{J}^{+}$ $=$ $-\sum_{\left\langle i,j\right\rangle}\sum_{\sigma=\uparrow,\downarrow}J_{\sigma}[\hat{n}_{i,\bar{\sigma}}\hat{c}_{i,\sigma}^{\dagger}\hat{c}_{j,\sigma}(1-\hat{n}_{j,\bar{\sigma}})$ (A2)
$+\hat{n}_{j,\bar{\sigma}}^{\dagger}\hat{c}_{j,\sigma}^{\dagger}\hat{c}_{i,\sigma}(1-\hat{n}_{i,\bar{\sigma}})],$
and $\hat{H}_{J}^{-}=(\hat{H}_{J}^{+})^{\dagger}$, in which $\bar{\sigma}$ denotes spin in the opposite direction of spin $\sigma$. While $\hat{H}_{J}^{0}$ describes the processes that do not change the number of doubly occupied sites, which has the form
$\hat{H}_{J}^{0}$ $=$ $-\sum_{\left\langle i,j\right\rangle}\sum_{\sigma=\uparrow,\downarrow}J_{\sigma}[\left(1-\hat{n}_{i,\bar{\sigma}}\right)\hat{c}_{i,\sigma}^{\dagger}\hat{c}_{j,\sigma}\left(1-\hat{n}_{j,\bar{\sigma}}\right)$ (A3)
$+\hat{n}_{i,\bar{\sigma}}\hat{c}_{i,\sigma}^{\dagger}\hat{c}_{j,\sigma}\hat{n}_{j,\bar{\sigma}}+\mathrm{H.c.}].$
The on-site and NN interaction terms are $\hat{H}_{U}=U\sum_{i}\hat{n}_{i,\uparrow}\hat{n}_{i,\downarrow}$ and $\hat{H}_{V}=(V/2)\sum_{\left\langle i,j\right\rangle}\hat{n}_{i}\hat{n}_{j}$, respectively.
Figure A1: Schematic illustrations of three types of hopping processes described by (a) $\hat{H}_{J}^{+}$, (b) $\hat{H}_{J}^{-}$ and (c) $\hat{H}_{J}^{0}$.
We assume that the on-site interaction is strong, that is $\left|U\right|\gg J_{\sigma},$ $\left|U\right|\gg\left|V\right|,$ and perform the Schrieffer-Wolff transformation
$\hat{H}_{\mathrm{eff}}$ $=$ $e^{\mathrm{i}\hat{S}}\hat{H}e^{-\mathrm{i}\hat{S}}$ (A4)
$=$ $\hat{H}+\mathrm{i}\left[\hat{S},\hat{H}\right]+\frac{\mathrm{i}^{2}}{2}\left[\hat{S},\left[\hat{S},\hat{H}\right]\right]+...$
$=$ $\hat{H}_{U}+\hat{H}_{V}+\hat{H}_{J}^{0}+\hat{H}_{J}^{+}+\hat{H}_{J}^{-}+\mathrm{i}\left[\hat{S},\hat{H}_{U}\right]$
$+\mathrm{i}\left[\hat{S},\hat{H}_{V}\right]+\mathrm{i}\left[\hat{S},\hat{H}_{J}^{+}+\hat{H}_{J}^{-}\right]+\mathrm{i}\left[\hat{S},\hat{H}_{J}^{0}\right]$
$+\frac{\mathrm{i}^{2}}{2}\left[\hat{S},\left[\hat{S},\hat{H}\right]\right]+...,$
where $\hat{S}$ requires to be a Hermitian operator, so that the above transformation is unitary. We firstly wish to eliminate the term $\hat{H}_{J}^{+}+\hat{H}_{J}^{-},$ which can be done by selecting the $\hat{S}$ operator such that $\hat{H}_{J}^{+}+\hat{H}_{J}^{-}+\mathrm{i}\left[\hat{S},\hat{H}_{U}\right]=0.$ Note that $\left[\hat{H}_{J}^{\pm},\hat{H}_{U}\right]=\mp U\hat{H}_{J}^{\pm}$, which suggests the form
$\hat{S}=-\frac{\mathrm{i}}{U}\left(\hat{H}_{J}^{+}-\hat{H}_{J}^{-}\right).$ (A5)
Therefore, we obtain
$\hat{H}_{\mathrm{eff}}$ $=$ $\hat{H}_{U}+\hat{H}_{V}+\hat{H}_{J}^{0}+\mathrm{i}\left[\hat{S},\hat{H}_{V}\right]+\frac{1}{U}\left[\hat{H}_{J}^{+},\hat{H}_{J}^{-}\right]$ (A6)
$+\mathrm{i}\left[\hat{S},\hat{H}_{J}^{0}\right]+O\left(J^{2}V/U^{2}\right)+O\left(J^{3}/U^{2}\right).$
Here we set $O(J_{\uparrow})=O(J_{\downarrow})=O(J).$ To eliminate the term $\left[\hat{S},\hat{H}_{J}^{0}\right],$ we note that $\hat{S}=O\left(J/U\right),$ thus we can consider a second order correction of $\hat{S}$ , that is
$\hat{S}\rightarrow\hat{S}+\hat{S}^{(2)},$ (A7)
with $\hat{S}^{(2)}=O\left(J^{2}/U^{2}\right)$ being determined by the operator equation $\mathrm{i}\left[\hat{S}^{(2)},\hat{H}_{U}\right]+\left[\hat{H}_{J}^{+}-\hat{H}_{J}^{-},\hat{H}_{J}^{0}\right]/U=0.$
Due to a large $\left|U\right|$, the doublons states are well separated from others with the energy gaps of order $\left|U\right|$. Then the Hamiltonian can be safely projected into the doublons subspace up to order $J^{2}/U$. The effective Hamiltonian in Eq. (A6) can be simplified by taking the following consideration into account: (i) Term $[\hat{S},\hat{H}_{V}]$ can be omitted, since it connects the doublons subspace to others. (ii) We neglect term $[\hat{S},[\hat{S},\hat{H}_{V}]]=O\left(J^{2}V/U^{2}\right)$, since it is higher than order $J^{2}/U$ and involves multi-site processes. (iii) Term $\hat{H}_{J}^{0}$ describing the hopping processes between singly occupied sites and empty (or doubly occupied) sites, also vanishes in the doublons subspace.
We finally need to evaluate the commutator $\left[\hat{H}_{J}^{+},\hat{H}_{J}^{-}\right].$ Since $\hat{H}_{J}^{+}$ and $\hat{H}_{J}^{-}$ are sums of pair operators, nonvanishing contributions arise only if the two pairs have one or two sites in common. Denoting $\hat{H}_{J}^{\pm}=\sum_{\left\langle i,j\right\rangle}\hat{H}_{J,(i,j)}^{\pm}$, we have
$\left[\hat{H}_{J}^{+},\hat{H}_{J}^{-}\right]=\sum_{\left\langle i,j\right\rangle}\sum_{\left\langle m,n\right\rangle}\left[\hat{H}_{J,(i,j)}^{+},\hat{H}_{J,(m,n)}^{-}\right]$
$=$ $\sum_{\left\langle i,j\right\rangle}\left[\hat{H}_{J,(i,j)}^{+},\hat{H}_{J,(i,j)}^{-}\right]+\sum_{\left\langle i,j,m\right\rangle}\left[\hat{H}_{J,(i,j)}^{+},\hat{H}_{J,(j,m)}^{-}\right],$
where $\sum_{\left\langle i,j,m\right\rangle}$ denotes the summation over trimers with site $j$ being a nearest neighbor of both sites $i$ and $m$. It can be checked that this three-site processes vanish in the doublons subspace, and the two-site processes remain. Direct calculation shows that
$\sum_{\left\langle i,j\right\rangle}\left[\hat{H}_{J,(i,j)}^{+},\hat{H}_{J,(i,j)}^{-}\right]$ (A9)
$=$ $(J_{\uparrow}^{2}+J_{\downarrow}^{2})\left[2\sum_{i}\hat{n}_{i,\uparrow}\hat{n}_{i,\downarrow}-\sum_{\left\langle i,j\right\rangle}(\hat{n}_{i,\uparrow}\hat{n}_{j,\downarrow}+\hat{n}_{i,\downarrow}\hat{n}_{j,\uparrow})\right]$
$+2J_{\uparrow}J_{\downarrow}\sum_{\left\langle i,j\right\rangle}(\hat{d}_{i}^{\dagger}\hat{d}_{j}+\hat{d}_{j}^{\dagger}\hat{d}_{i}+\hat{s}_{i}^{+}\hat{s}_{j}^{-}+\hat{s}_{j}^{+}\hat{s}_{i}^{-}),$
where $\hat{s}_{i}^{+}=\hat{c}_{i,\uparrow}^{\dagger}\hat{c}_{i,\downarrow}$, $\hat{s}_{i}^{-}=(\hat{s}_{i}^{+})^{\dagger}$ are the spin-flip operator, and $d_{i}^{\dagger}=\hat{c}_{i,\uparrow}^{\dagger}\hat{c}_{i,\downarrow}^{\dagger}$ is the creation operator of doublon. Note that in the doublons subspace, terms $\hat{s}_{i}^{+}\hat{s}_{j}^{-}$ and $\hat{s}_{i}^{z}\hat{s}_{j}^{z}=(\hat{n}_{i,\uparrow}-\hat{n}_{i,\downarrow})(\hat{n}_{j,\uparrow}-\hat{n}_{j,\downarrow})/4$ vanish, and the on-site interacting term gives constant energy for a fixed number of doublons. Then term $-2\hat{s}_{i}^{z}\hat{s}_{j}^{z}$ can be added to Eq. (A9). Up to order $J^{2}/U$, the effective Hamiltonian that conserves the total number of doublons is
$\hat{H}_{\mathrm{eff}}=\sum_{\left\langle i,j\right\rangle}\left[J_{\mathrm{eff}}\left(\hat{d}_{i}^{\dagger}\hat{d}_{j}+\hat{d}_{j}^{\dagger}\hat{d}_{i}\right)+\frac{1}{2}V_{\mathrm{eff}}\hat{n}_{i}\hat{n}_{j}\right],$ (A10)
where the effective hopping amplitude and NN interaction of doublons are $J_{\mathrm{eff}}=2J_{\uparrow}J_{\downarrow}/U$ and $V_{\mathrm{eff}}=V-(J_{\uparrow}^{2}+J_{\downarrow}^{2})/U$, respectively, and the constant term is omitted.